# Add-One-In: Incremental Sample Selection for Large Language Models via a Choice-Based Greedy Paradigm

### v1

@@ 4.3 Scaling-Up with Larger Datasets
We conduct experiments in the medical domain to demonstrate the practical effectiveness of our method. Specifically, based on the HuatuoGPT-sft-data-v1 dataset for SFT samples, we evaluate performance on the Chinese Medical Benchmark (CMB) Wang et al. (2024), utilizing the Baichuan2-7B-Chat model Baichuan (2023). Given that AlpaGasus requires querying ChatGPT for sample scoring, we primarily compare our method with the Base model, Full, and IFD. We provide detailed training settings in App. B.6.
@@ 6 Limitations and Future Work
While our proposed choice-based sample selection paradigm demonstrates promising results, it is important to acknowledge several limitations and potential areas for future improvement. A significant limitation of our approach is its reliance on LLMs whose capabilities and biases directly impact the sample selection process. The rationality and effectiveness of selecting samples are contingent upon the LLM’s ability to accurately assess and compare data points. However, LLMs may have inherent biases inherited from their training data, which can inadvertently influence the selection process Wang et al. (2023); Ko et al. (2020). Another limitation lies in the need to manually adjust hyperparameters such as the sizes of sets $\mathcal{A}^{\prime}$ and $\mathcal{B}^{\prime}$, as well as the design of prompts used to elicit evaluations from the LLM. Selecting the sizes of $\mathcal{A}^{\prime}$ and $\mathcal{B}^{\prime}$ impacts the granularity, and breadth of the selection process, while the choice of prompts influences the quality of the LLM’s assessments.
In the future, enhancing the greedy sampling process by incorporating more sophisticated heuristic methods instead of random selection offers a promising direction. Besides, how to extend my approach to online and sequential data scenarios remains an open challenge. For example, developing adaptive sampling strategies would significantly enhance the practicality of our method. This could involve designing algorithms capable of making immediate selection decisions in real-time applications like NER Hu et al. (2023) and Summarization Hu et al. (2022), ensuring sustained model performance in environments where data arrives continuously.
Appendix A Algorithm
Input: $\mathcal{D}$: Full dataset, a LLMs and hyper-parameters: {$K$: Desired subset size, $L_{A}$ and $L_{B}$: window size.}
Output: $\mathcal{A}$: Selected subset
3
4while $|\mathcal{A}|<K$ do
6
8       update $\mathcal{A}\leftarrow\mathcal{A}\cup\{x^{*}\}$;
9       update $\mathcal{B}\leftarrow\mathcal{B}\setminus\{x^{*}\}$;
10
return $\mathcal{A}$
Algorithm 1 Algorithm Process of Our Method.
Appendix B Experiment
B.1 Implementation Details and Cost
@@
B.2 Pairwise Comparison with Random Baseline
As depicted in Fig.7, our method consistently outperforms the Random baseline across various settings by significant margins, demonstrating the excellent effectiveness of the proposed method.
B.3 Ablation Study
@@
Intuitively, a larger $L_{A}$ and $L_{B}$ results in $\mathcal{A}^{\prime}$ and $\mathcal{B}^{\prime}$ containing more elements, which increases the number of candidates and simultaneously elevates the difficulty of assessing diversity among them. This expansion in the candidate subsets provides the LLM with more options to consider, potentially leading to the selection of samples with greater diversity. Our experimental results are summarized in Tab. 6, where we compare different configurations in terms of their ability to select better samples. We report the number of times one configuration wins over another, indicating superiority in diversity and quality of the selected subsets. We conduct evaluations using various LLMs as selectors (Qwen2.5-72B-Instruct and ChatGPT) on multiple datasets (e.g., Sinstruct, Vicuna, Koala, WizardLM). From the results in Tab. 6, we observe several key trends:
@@
  2. 2.
    Effect of Length of Candidate Set $\mathcal{B}^{\prime}$: Similarly, a larger candidate set $\mathcal{B}^{\prime}$ (length 20) improves performance compared to a smaller candidate set (length 10). For example, on the WizardLM dataset with ChatGPT as the selector, $\mathcal{B}^{\prime}$ of length 20 wins 85 times versus 68 wins for length 10. A larger candidate pool offers more options for the LLM to select diverse and high-quality samples.
@@
  5. 5.
    Consistency Across Models and Datasets: These trends are observed across different LLM selectors (Qwen2.5-72B-Instruct and ChatGPT) and datasets (Sinstruct, WizardLM, Vicuna, Koala). This consistency suggests that the benefits of larger set sizes and multi-round selection are generalizable. Besides, our mehto is not limited to specific models or datasets.
@@
B.4 Explanation To Characteristics Metrics
@@
TTR=Number of unique typesNumber of $\text{TTR}=\frac{\text{Number of unique types}}{\text{Number of tokens}}*100$ (3)
@@
MTLD=Total number of words in the textNumber of $\text{MTLD}=\frac{\text{Total number of words in the text}}{\text{Number of factors}}$ (4)
@@
B.5 In-Depth Analysis of Efficiency
@@
Beyond theoretical analysis, we provide empirical comparisons of the actual time consumed during the selection process. Based on the Alpaca dataset and under identical hardware conditions, we reproduce IFD and Alpagasus for fair comparison. For both IFD and our method, we utilize Qwen2.5-7B-Instruct and set $L_{A}=L_{B}=20$ in our method. For Alpagasus, we use the official OpenAI API to query ChatGPT. To conserve time and computational resources, we measure the time required to perform selection (for our method) or scoring (for IFD and Alpagasus) on 1,000 samples. We report the total time taken for the process.
Additionally, we also explore how varying the value of $L$ in our method affects the overall practical time consumption (PTC), providing insights into the trade-offs between selection granularity and computational efficiency. Specifically, under the same hardware environment, we record the time consumption for 100 selections under different values of $L_{A}$ and $L_{B}$, when employing Llama3.1-8B-Instruct as the selector. he results in Tab. 7 show that our method does not experience a sharp increase in time consumption as $L_{A}$ and $L_{B}$ rise. Although the practical time consumption generally grows with larger $L_{A}$ and $L_{B}$ values, the rate of increase is relatively moderate. This indicates that within a certain range, we can adjust $L_{A}$ and $L_{B}$ to balance the selection and computational cost without being overly burdened by excessive time expenditure.
@@
B.6 Scaling-up with Larger Dataset
@@
We adopt HuatuoGPT-sft-data-v1333https://huggingface.co/datasets/FreedomIntelligence/HuatuoGPT-sft-data-v1 as SFT samples that includes 226K Chinese medical QA pairs as SFT samples and evaluate the performance on Chinese Medical Benchmark (CMB) Wang et al. (2024) that is a comprehensive and all-encompassing Chinese medical quiz assessment benchmark, which contains 12K human-annotated five-option multi-choice questions and cover six aspects in benchmarking a medical LLM.
Hyper-parameters
The Tab. 8 highlights the training hyper-parameters.
@@
Appendix C Case Analysis
In this section, we provide three selection cases analysis when employing ChatGPT as the selector. As shown in Fig. 8, Fig. 9, and Fig. 10, we find LLMs can effectively make a reasonable selection choice given the selection prompt that includes the selected set $A$ and the candidate set $B$. For example, as shown in Fig. 8, the selector recommend [A]-Element for the optimal element to add to the set $A$, by providing a highly informative response on the Quicksort algorithm and introduces a distinct topic to Set $A$. The other elements in Candidate Set $B$ also demonstrate high response quality but do not introduce as unique of a concept as the Quicksort algorithm in [A]-Element.

### v2

@@ 4.3 Scaling-Up with Larger Datasets
We conduct experiments in the medical domain to demonstrate the practical effectiveness of our method. Specifically, based on the HuatuoGPT-sft-data-v1 dataset for SFT samples, we evaluate performance on the Chinese Medical Benchmark (CMB) Wang et al. (2024), utilizing the Baichuan2-7B-Chat model Baichuan (2023). Given that AlpaGasus requires querying GPT3.5 for sample scoring, we primarily compare our method with the Base model, Full, and IFD. We provide detailed training settings in App. A.6.
@@ 5 Conclusion
6 Acknowledgments
This work was supported by the Guangxi Key R&D Project (No. AB24010167), the Project (No. 20232ABC03A25), and the Futian Healthcare Research Project (No.FTWS002). This work was also supported by the National Natural Science Foundation of China (NSFC) under Grant 62402158.
@@ 7 Limitations and Future Work
While our proposed choice-based sample selection paradigm demonstrates promising results, it is important to acknowledge several limitations and potential areas for future improvement. A significant limitation of our approach is its reliance on LLMs whose capabilities and biases directly impact the sample selection process. The rationality and effectiveness of selecting samples are contingent upon the LLM’s ability to accurately assess and compare data points. However, LLMs may have inherent biases inherited from their training data, which can inadvertently influence the selection process Wang et al. (2023); Ko et al. (2020). Another limitation lies in the need to manually adjust hyperparameters such as the sizes of sets $\mathcal{A}^{\prime}$ and $\mathcal{B}^{\prime}$, as well as the design of prompts used to elicit evaluations from the LLM. Selecting the sizes of $\mathcal{A}^{\prime}$ and $\mathcal{B}^{\prime}$ impacts the granularity and breadth of the selection process, while the choice of prompts influences the quality of the LLM’s assessments.
In the future, enhancing the greedy sampling process by incorporating more sophisticated heuristic methods instead of random selection offers a promising direction. Besides, how to extend my approach to online and sequential data scenarios remains an open challenge. For example, developing adaptive sampling strategies would significantly enhance the practicality of our method. This could involve designing algorithms capable of making immediate selection decisions in real-time applications, ensuring sustained model performance in environments where data arrives continuously.
Appendix A Experiment
A.1 Implementation Details and Cost
@@
A.2 Performance of Benchmark on Varying Training Samples
In this section, we conduct an ablation study to analyze the impact of varying training data proportions on the performance of LLMs across different benchmarks. Instead of ablating components of a model architecture, we systematically evaluate models trained on increasing subsets of the full training data: 3K, 6K, and 9K examples. This allows us to isolate the effect of training data size on the model’s ability to generalize to unseen tasks, as measured by MMLU, BBH, and Hellaswag. By comparing these results to the performance of a model trained on the complete dataset (52K), we aim to quantify the performance gains achieved with larger training sets and identify potential saturation points or diminishing returns. This analysis provides insights into the data efficiency of LLMs and the importance of training data scale for achieving optimal benchmark performance.
Table 7 presents a comprehensive performance comparison across several few-shot learning methods: Full (utilizing a larger 52K dataset), Random (subsampling to 3K, 6K, and 9K), Alpagasus (3K, 6K, and 9K), IFD (3K, 6K, and 9K), and our proposed approach, “Ours” (evaluated on 3K, 6K, and 9K subsets). The evaluation spans diverse and challenging benchmarks, including MMLU (assessed in both 0-shot and 5-shot settings), BBH, and Hellaswag. The results highlight the performance trends as the size of the training data increases from 3K to 9K, providing insights into the scalability and effectiveness of each method on different tasks. Notably, the final column for each data size (9K) emphasizes the performance of “OURS” often demonstrating competitive or superior results in these higher data regimes. The average performance across all tasks is also provided at the bottom, offering a holistic view of each method’s overall effectiveness.
Table 6: Performance Comparison
Benchmarks Full Random Alpagasus IFD OURS Random Alpagasus IFD OURS Random Alpagasus IFD OURS
(52K) (3K) (3K) (3K) (3K) (6K) (6K) (6K) (6K) (9K) (9K) (9K) (9K)
MMLU-0-Shot 22.09 21.10 22.00 21.72 22.30 21.80 22.40 22.10 22.70 22.51 23.25 23.12 23.82
MMLU-5-Shot 45.45 44.00 44.70 44.80 45.60 45.00 45.70 45.30 46.10 46.38 46.74 46.10 47.44
BBH 32.10 30.50 30.80 31.40 31.20 30.90 31.10 31.60 31.50 31.38 31.32 31.08 30.97
Hellaswag 69.97 69.00 69.80 69.30 70.10 69.50 71.20 69.65 70.50 70.99 71.07 70.55 71.07
Average 42.15 41.15 41.83 41.53 42.30 41.80 42.60 41.97 42.70 42.82 43.10 42.71 43.33
@@ A.3 Ablation Study
Intuitively, a larger $L_{A}$ and $L_{B}$ results in $\mathcal{A}^{\prime}$ and $\mathcal{B}^{\prime}$ containing more elements, which increases the number of candidates and simultaneously elevates the difficulty of assessing diversity among them. This expansion in the candidate subsets provides the LLM with more options to consider, potentially leading to the selection of samples with greater diversity. Our experimental results are summarized in Tab. 8, where we compare different configurations in terms of their ability to select better samples. We report the number of times one configuration wins over another, indicating superiority in diversity and quality of the selected subsets. We conduct evaluations using various LLMs as selectors (Qwen2.5-72B-Instruct and GPT3.5) on multiple datasets (e.g., Sinstruct, Vicuna, Koala, WizardLM). From the results in Tab. 8, we observe several key trends:
@@
    Effect of the initialization of $\mathcal{A}^{\prime}$: In the preceding experiment, the $\mathcal{A}^{\prime}$ was initialized through random sampling. It is hypothesized that a more guided initialization, such as employing K-means clustering, could lead to a more advantageous final subset selection and consequently improved model performance. In this section, we experimentally investigate the impact of various initialization strategies for $\mathcal{A}^{\prime}$ on the characteristics of the ultimately chosen subset. The results presented in Table 9 indicate that utilizing K-means and KNN for the initialization of $\mathcal{A}^{\prime}$ does indeed result in slightly enhanced Diversity Measurements, as evidenced by metrics like TTR and MTLD. However, when considering Quality Measurement, the Random Initialization approach achieves comparable performance to both K-means and KNN, even outperforming them in terms of Helpful Score. These findings suggest that our method exhibits a degree of robustness with respect to different initializations of $\mathcal{A}^{\prime}$. We posit that the underlying reason for this resilience stems from our experimental constraint of setting the size of $\mathcal{A}^{\prime}$ to a maximum of 20, as a consequence, the initial selection within a relatively small pool might not drastically alter the final refined subset obtained through our subsequent selection process.
  2. 2.
@@
    Effect of Length of Candidate Set $\mathcal{B}^{\prime}$: Similarly, a larger candidate set $\mathcal{B}^{\prime}$ (length 20) improves performance compared to a smaller candidate set (length 10). For example, on the WizardLM dataset with GPT3.5 as the selector, $\mathcal{B}^{\prime}$ of length 20 wins 85 times versus 68 wins for length 10. A larger candidate pool offers more options for the LLM to select diverse and high-quality samples.
  4. 4.
@@
  6. 6.
    Consistency Across Models and Datasets: These trends are observed across different LLM selectors (Qwen2.5-72B-Instruct and GPT3.5) and datasets (Sinstruct, WizardLM, Vicuna, Koala). This consistency suggests that the benefits of larger set sizes and multi-round selection are generalizable. Besides, our method is not limited to specific models or datasets.
Method Diversity Quality #Tokens
Full (52K) 95.47 7.9352 0.1067 2.765 1.314 11.35
Random 95.46 7.9421 0.1068 2.764 1.397 11.33
AlpaGasus 96.07 8.0508 0.1086 2.969 2.025 10.93
IFD 96.05 8.0444 0.1091 3.127 2.456 10.80
OURS + Random 96.24 8.4449 0.1035 3.210 2.703 11.29
OURS + Kmeans 96.58 9.0125 0.1057 3.092 2.613 11.57
OURS + KNN 96.39 8.8721 0.1048 3.278 2.581 11.98
@@
A.4 Explanation To Characteristics Metrics
@@
$\text{TTR}=\frac{\text{Number of unique types}}{\text{Number of tokens}}*100$ (3)
@@
[…]
@@
A.5 In-Depth Analysis of Efficiency
@@
Beyond theoretical analysis, we provide empirical comparisons of the actual time consumed during the selection process. Based on the Alpaca dataset and under identical hardware conditions, we reproduce IFD and Alpagasus for fair comparison. For both IFD and our method, we utilize Qwen2.5-7B-Instruct and set $L_{A}=L_{B}=20$ in our method. For Alpagasus, we use the official OpenAI API to query GPT3.5. To conserve time and computational resources, we measure the time required to perform selection (for our method) or scoring (for IFD and Alpagasus) on 1,000 samples. We report the total time taken for the process.
Additionally, we also explore how varying the value of $L$ in our method affects the overall practical time consumption (PTC), providing insights into the trade-offs between selection granularity and computational efficiency. Specifically, under the same hardware environment, we record the time consumption for 100 selections under different values of $L_{A}$ and $L_{B}$, when employing Llama3.1-8B-Instruct as the selector. The results in Tab. 10 show that our method does not experience a sharp increase in time consumption as $L_{A}$ and $L_{B}$ rise. Although the practical time consumption generally grows with larger $L_{A}$ and $L_{B}$ values, the rate of increase is relatively moderate. This indicates that within a certain range, we can adjust $L_{A}$ and $L_{B}$ to balance the selection and computational cost without being overly burdened by excessive time expenditure.
@@
A.6 Scaling-up with Larger Dataset
@@
We adopt HuatuoGPT-sft-data-v1222https://huggingface.co/datasets/FreedomIntelligence/HuatuoGPT-sft-data-v1 as SFT samples that includes 226K Chinese medical QA pairs as SFT samples and evaluate the performance on Chinese Medical Benchmark (CMB) Wang et al. (2024) that is a comprehensive and all-encompassing Chinese medical quiz assessment benchmark, which contains 12K human-annotated five-option multi-choice questions and cover six aspects in benchmarking a medical LLM. The Tab. 11 highlights the training hyper-parameters.
@@
Appendix B Case Analysis
In this section, we provide three case selection analyses when employing GPT3.5 as the selector. As shown in Fig. 8, Fig. 9, and Fig. 10, we find LLMs can effectively make a reasonable selection choice given the selection prompt that includes the selected set $A$ and the candidate set $B$. For example, as shown in Fig. 8, the selector recommends [A]-Element for the optimal element to add to the set $A$, by providing a highly informative response on the Quicksort algorithm and introducing a distinct topic to Set $A$. The other elements in Candidate Set $B$ also demonstrate high response quality but do not introduce as unique of a concept as the Quicksort algorithm in [A]-Element.